# Axis Decomposition for ODRL: Resolving Dimensional Ambiguity in Policy Constraints through Interval Semantics

### v1

@@ Abstract
Every ODRL 2.2 constraint compares a single scalar value: (leftOperand, operator, rightOperand). Five of ODRL’s  34 left operands, however, denote multi-dimensional quantities—image dimensions, canvas positions, geographic coordinates—whose specification text explicitly references multiple axes. For these operands, a single scalar constraint admits one interpretation per axis, making policy evaluation non-deterministic. We classify ODRL’s left operands by value-domain structure (scalar, dimensional, concept-valued), grounded in the ODRL 2.2 specification text, and show that dimensional ambiguity is intrinsic to the constraint syntax. We present an axis-decomposition framework that refines each dimensional operand into axis-specific scalar operands and prove four properties: deterministic interpretation, AABB completeness, sound over-approximation under projection, and conservative extension. Conflict detection operates in two layers: per-axis verdicts are always decidable; box-level verdicts compose through Strong Kleene conjunction into a three-valued logic (Conflict, Compatible, Unknown). For ODRL’s disjunctive (odrl:or) and exclusive-or (odrl:xone) logical constraints, where per-axis decomposition does not apply, the framework encodes coupled multi-axis conjectures directly. We instantiate the framework as the ODRL Spatial Axis Profile—15 axis-specific left operands for the five affected base terms—and evaluate it on 117 benchmark problems spanning nine categories across both TPTP FOF (Vampire) and SMT-LIB (Z3) encodings, achieving full concordance between provers. Benchmark scenarios are inspired by constraints arising in cultural heritage dataspaces such as Datenraum Kultur. All meta-theorems are mechanically verified in Isabelle/HOL.
@@ 1 Introduction
ODRL Spatial Axis Profile: 15 axis-specific left operands, OWL vocabulary, SHACL validation, and TPTP axiom files (Section˜3.6).
117 dual-encoded benchmarks (TPTP FOF + SMT-LIB) in nine categories, with 100% Vampire/Z3 concordance (Section˜3.7).
@@ 3.2.2 Box-Level Verdicts
Table 3: Verdict correspondence across reasoning layers. The ODRL verdict maps to SZS ontology status (TPTP/Vampire) and SMT-LIB status (Z3) depending on problem encoding.
ODRL Verdict Encoding SZS Status Z3 Status
@@
Compatible $\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket\stackrel{{\scriptstyle?}}{{\neq}}\emptyset$ Theorem sat
Unknown (axis unconstrained) GaveUp / n/a unknown
@@ 3.4 Policy Quality Analysis
The exact domain $D_{k}$ for each axis-specific operand is fixed by the profile instantiation (Section˜3.6). For size operands, $D_{k}=(0,\infty)$ since zero collapses the dimension; for relative positions, $D_{k}=[0,100]$; for geographic coordinates, $D_{k}$ follows WGS 84 bounds. All theorems in this section are parametric in $D_{k}$ and hold for any totally ordered domain—open, closed, bounded, or unbounded.
@@ 3.5 Mechanical Verification
Verified results: Theorems 3.33.19, Lemma 3.12, and end-to-end bridge theorems connecting verdict-level results to set-theoretic properties. The BSB scenario is included as a concrete regression test. The full proof script builds in under 30 s with no sorry. [^2] Isabelle theories and benchmark suite available upon publication.
3.6 Profile Implementation
We instantiate the axis-decomposition framework as the ODRL Spatial Axis Profile (OAX), consisting of three artefacts:
An OWL 2 ontology at http://w3id.org/odrl/spatial-axis# declares 15 axis-specific LeftOperand instances—three per dimensional base term—with SKOS definitions linking each axis operand to its parent ODRL operand and specifying the value domain.
A SHACL shapes graph enforces domain-bound validation at policy authoring time. Each axis operand has a corresponding sh:NodeShape that constrains the odrl:rightOperand to xsd:decimal with domain-appropriate bounds:
  • size axes (absoluteSizeWidth, …): $\mathtt{minExclusive}\;0$, matching domain $(0,\infty)$;
  • geographic axes: [90,90][-90,90] for latitude, [180,180][-180,180] for longitude;
  • relative position axes: $[0,100]$ (percentage).
These shapes serve a dual role: (i) they catch malformed policies before reasoning (e.g. latitude =200=200), and (ii) they ground the domain bounds that appear as axioms in the TPTP encoding.

### v2

@@ Abstract
Every ODRL 2.2 constraint compares a single scalar value: (leftOperand, operator, rightOperand). Five of ODRL’s left operands, however, denote multi-dimensional quantities—image dimensions, canvas positions, geographic coordinates—whose specification text explicitly references multiple axes. For these operands, a single scalar constraint admits one interpretation per axis, making policy evaluation non-deterministic. We classify ODRL’s left operands by value-domain structure (scalar, dimensional, concept-valued), grounded in the ODRL 2.2 specification text, and show that dimensional ambiguity is intrinsic to the constraint syntax. We present an axis-decomposition framework that refines each dimensional operand into axis-specific scalar operands and prove four properties: deterministic interpretation, AABB completeness, sound over-approximation under projection, and conservative extension. Conflict detection operates in two layers: per-axis verdicts are always decidable; box-level verdicts compose through Strong Kleene conjunction into a three-valued logic (Conflict, Compatible, Unknown). For ODRL’s disjunctive (odrl:or) and exclusive-or (odrl:xone) logical constraints, where per-axis decomposition does not apply, the framework encodes coupled multi-axis conjectures directly. We instantiate the framework as the ODRL Spatial Axis Profile—15 axis-specific left operands for the five affected base terms—and evaluate it on 117 benchmark problems spanning nine categories across both TPTP FOF (Vampire) and SMT-LIB (Z3) encodings, achieving full concordance between provers. Benchmark scenarios are inspired by constraints arising in cultural heritage dataspaces such as Datenraum Kultur. All meta-theorems are mechanically verified in Isabelle/HOL.
@@ 1 Introduction
ODRL Spatial Axis Profile: 15 axis-specific left operands, OWL vocabulary, SHACL validation, and TPTP axiom files (Footnote˜2).
117 dual-encoded benchmarks (TPTP FOF + SMT-LIB) in nine categories, with 100% Vampire/Z3 concordance (Section˜3.6).
@@ 3.2.2 Box-Level Verdicts
Table 3: Verdict correspondence across reasoning layers. Per-axis verdicts map to prover outputs; Unknown arises at the box level when an axis lacks constraints and is not submitted to the prover.
Relation Verdict Encoding SZS Z3
Conflict Conflict $\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket=\emptyset$ Theorem unsat
Compatible $\llbracket c_{1}\rrbracket\cap\llbracket c_{2}\rrbracket\neq\emptyset$ Theorem sat
@@
Refuted $\llbracket c_{1}\rrbracket\not\subseteq\llbracket c_{2}\rrbracket$ CounterSat sat
Unknown (either relation) (axis unconstrained) — not submitted —
[…]
@@
Definition 3.17 (Request Satisfaction).
For request $\mathbf{v}=(v_{1},\ldots,v_{n})$ and constraint set $C$:
$\operatorname{satisfied}(\mathbf{v},C)=\begin{cases}\textsc{Yes}&\text{if }\forall k:v_{k}\in I_{k}\\ \textsc{No}&\text{if }\exists k:v_{k}\notin I_{k}\end{cases}$
Remark 3.18 (Deontic Conflict).
When comparing a permission and prohibition over the same action, deontic conflict occurs when scopes overlap: $\llbracket C_{P}\rrbracket\cap\llbracket C_{Q}\rrbracket\neq\emptyset$. This is the inverse of constraint satisfiability (Definition 3.10): disjoint scopes prevent deontic conflict, while overlapping scopes create it.
@@ 3.4 Policy Quality Analysis
The exact domain $D_{k}$ for each axis-specific operand is fixed by the profile instantiation (Footnote˜2). For size operands, $D_{k}=(0,\infty)$ since zero collapses the dimension; for relative positions, $D_{k}=[0,100]$; for geographic coordinates, $D_{k}$ follows WGS 84 bounds. All theorems in this section are parametric in $D_{k}$ and hold for any totally ordered domain—open, closed, bounded, or unbounded.
@@ 3.5 Mechanical Verification
Verified results: Theorems 3.33.21, Lemma 3.12, and end-to-end bridge theorems connecting verdict-level results to set-theoretic properties. The BSB scenario is included as a concrete regression test. The full proof script builds in under 30 s with no sorry. [^2] Isabelle theories and benchmark suite available upon publication. We instantiate the framework as the ODRL Spatial Axis Profile (OAX): an OWL 2 ontology (http://w3id.org/odrl/spatial-axis#) declaring 15 axis-specific LeftOperand instances with SKOS links to parent operands; a SHACL shapes graph constraining odrl:rightOperand to xsd:decimal with domain-appropriate bounds (size: $(0,\infty)$; latitude: [90,90][-90,90]; longitude: [180,180][-180,180]; relative: $[0,100]$); and TPTP axiom files encoding the interval predicates. The SHACL shapes catch malformed policies at authoring time and ground the domain bounds used in automated reasoning.